OpenStudy (adunb8):

this questions is for differential equation & linear algebra level. can anyone tell me about separable equations: ex: yy'=4x, y(1) = -3

OpenStudy (anonymous):

ydy = 4xdx Since it already separate, take integral both sides

OpenStudy (adunb8):

oh if it is separable i have to use v=y/x which y= xv and y' = v+xv'

OpenStudy (adunb8):

and how do you know if its separate or not?

OpenStudy (anonymous):

Separate means x terms in one side and y terms in the other side!

OpenStudy (adunb8):

what if its like 3y^2y' = (1+y^3)cosx like this problem i wouldnt know right away that it is separable or not how do i tell ?

OpenStudy (anonymous):

Because ( 1 + y^3) to the right, you need to switch them to the left side!

OpenStudy (adunb8):

so all i am doing for separable equation is to separate x to one side and y to other?

OpenStudy (anonymous):

ydy = 4xdx Since it already separate, take integral both sides

OpenStudy (adunb8):

oh i see okay so what would be the total solution?

OpenStudy (anonymous):

=> y = ..

OpenStudy (adunb8):

?? y = ...?